In a thermocouple thermometer connected to a millivoltmeter, which can read up to 10mv. One junction is in ice at 0°c and the other in steam at 100°c, the millimeter reads 4.0mv. What is the maximum temperature, which the arrangement can measure. 

Grade:10

1 Answers

Susmita
5 years ago
In a thermocouple,emf
E=at+bt^2
t is temperature difference between cold and hot junctions.
Here no information about a,b is given.So we have to take the linear term only .
E=at and keep the cold reservoir fixed at 0 degree centigrade.
Given E=4 mV when t=100°C.
So a=E/t=0.04 mV/°C.
Maximum value of t will correspond to maximum emf or E=10mV.
So t=E/a=10/.04=250°C.
Please approve if helped.